OpenStudy (anonymous):

hi how do you do 3:4=18:-----

OpenStudy (anonymous):

18/3 = 6 so ,multiply by 6 to 4 that will be ur answer

OpenStudy (anonymous):

or you can write 3/4 =18/x so x= 18*4/3

OpenStudy (anonymous):

which is 4*6 = 24

OpenStudy (anonymous):

but how do you remember it for a test

OpenStudy (anonymous):

okay you remember this a:b = c:d can be written as \[\frac{ a }{ b }=\frac{ c }{ d }\]

OpenStudy (anonymous):

is that the formula for ratio

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

yaw

OpenStudy (anonymous):

Is rates same aa ratios

OpenStudy (anonymous):

I find it harder

OpenStudy (anonymous):

rates as in?

ganeshie8 (ganeshie8):

\(\large a : b\) is same as \(\large \dfrac{a}{b}\)

OpenStudy (anonymous):

i think in that u just have to see by how much was it increased or decreased

OpenStudy (anonymous):

calculatig rates, writing rates, and converting rates

ganeshie8 (ganeshie8):

yes they all are the problems which use `ratios`

OpenStudy (anonymous):

yeah,similar to it

ganeshie8 (ganeshie8):

ratio is mathematical construct like `division` or `addition` it is used in problems, like u use addition/subtraction etc...

ganeshie8 (ganeshie8):

do u have any example of converting rate problem ?

OpenStudy (anonymous):

5 apples bought for 60c is a rate of_______ c per apple/

ganeshie8 (ganeshie8):

ahh ok :) 5apples : 60c = 1apple : ____

ganeshie8 (ganeshie8):

\(\large \dfrac{5}{60} = \dfrac{1}{x}\) solve \(x\)

ganeshie8 (ganeshie8):

indeed its a very good application of ratios !

OpenStudy (anonymous):

ahah thankyou